Refer to the individual rate schedules in Appendix C. a. What are the tax liability, the marginal tax rate, and the average tax rate for a married couple filing jointly with $64,900 taxable income? b. What are the tax liability, the marginal tax rate, and the average tax rate for a single individual with $193,300 taxable income? c. What are the tax liability, the marginal tax rate, and the average tax rate for a head of household with $459,300 taxable income? Complete this question by entering your answers in the tabs below. Required A Required B Required C What are the tax liability, the marginal tax rate, and the average tax rate for a married couple filing jointly with $64,900 taxable income? (Round final Tax liability" to the nearest whole dollar and "Average tax rate" to 2 decimal places.) Tax liability Marginal tax rate Average tax rate Required B Required A
Complete this question by entering your answers in the tabs below. Required A Required B Required C What are the tax liability, the marginal tax rate, and the average tax rate for a single individual with $193,300 taxable income? (Do not round your intermediate computations. Round final "Tax liability" to the nearest whole dollar and "Average tax rate" to 2 decimal places.) Tax liability % Marginal tax rate Average tax rate % Required A Required C
Complete this question by entering your answers in the tabs below. Required C Required A Required B What are the tax liability, the marginal tax rate, and the average tax rate for a head of household with $459,300 taxable income? (Round final "Tax liability" to the nearest whole dollar and "Average tax rate" to 2 decimal places.) Tax liability Marginal tax rate % Average tax rate % Required B Required C
Answer #1
a) Tax liability = 1905+(64900-19050)*12% = $   7,407.00
Marginal tax rate 12.00%
[Marginal tax rate is the rate applicable to the next
1$ of taxable income. The rate is 12% upto $77400]
Average tax rate = 7407/64900 = 11.41%
b) Tax liability = 32089.50+(193300-157500)*32% = $       43,546
Marginal tax rate 32.00%
Average tax rate = 43546/193300 = 22.53%
c) Tax liability = 44298+(459300-200000)*35% = $   1,35,053
Marginal tax rate 35.00%
Average tax rate = 135053/459300 = 29.40%